# PBEBench: A Multi-Step Programming by Examples Reasoning Benchmark inspired by Historical Linguistics

### v3

@@ G.1 Effect of More Examples
G.2 Efficiency of Problem Proposer
@@
G.3 Factorial Analysis
@@
G.3.1 Effect of Long Chain of Thought Reasoning
@@
G.3.2 Analysis of Models of Varying Capabilities
@@
G.4 Logistic Regression Analysis
@@
Cascade
Length
Pass@1
@@
Sampling
Budget
First Code Block Last Code Block Nulls
@@
Sampling
Budget
First Code Block Last Code Block Nulls
@@
Max Seq
Length
First Code Block Last Code Block Nulls
@@
G.5 Confusion Matrices for Cascade Lengths
@@
G.6 Confusion Matrices for BFCC Relations
@@
Figure 29: Relation Type Confusion Matrix for Failure (Claude-3.5-Sonnet PBEBench-Lite): Confusion matrix showing the distribution of relation types in the model prediction vs ground truth. INVALID category indicates cases where the model fails to generate a valid cascade at all. The results are for Claude-3.5-Sonnet on PBEBench-Lite, and failure denotes Pass@1 = 0 with a sampling budget of 1.
Figure 30: Relation Type Confusion Matrix for Success (Claude-3.5-Sonnet PBEBench-Lite): Confusion matrix showing the distribution of relation types in the model prediction vs ground truth. INVALID category indicates cases where the model fails to generate a valid cascade at all. The results are for Claude-3.5-Sonnet on PBEBench-Lite, and success denotes Pass@1 = 1 with a sampling budget of 1.
Figure 31: Relation Type Confusion Matrix for Failure (Claude-3.7-Sonnet (Thinking) PBEBench-Lite): Confusion matrix showing the distribution of relation types in the model prediction vs ground truth. INVALID category indicates cases where the model fails to generate a valid cascade at all. The results are for Claude-3.7-Sonnet (Thinking) on PBEBench-Lite, and failure denotes Pass@1 = 0 with a sampling budget of 1.
Figure 32: Relation Type Confusion Matrix for Success (Claude-3.7-Sonnet (Thinking) PBEBench-Lite): Confusion matrix showing the distribution of relation types in the model prediction vs ground truth. INVALID category indicates cases where the model fails to generate a valid cascade at all. The results are for Claude-3.7-Sonnet (Thinking) on PBEBench-Lite, and success denotes Pass@1 = 1 with a sampling budget of 1.
Figure 33: Relation Type Confusion Matrix for Failure (Claude-3.7-Sonnet PBEBench-Lite): Confusion matrix showing the distribution of relation types in the model prediction vs ground truth. INVALID category indicates cases where the model fails to generate a valid cascade at all. The results are for Claude-3.7-Sonnet on PBEBench-Lite, and failure denotes Pass@1 = 0 with a sampling budget of 1.
Figure 34: Relation Type Confusion Matrix for Success (Claude-3.7-Sonnet PBEBench-Lite): Confusion matrix showing the distribution of relation types in the model prediction vs ground truth. INVALID category indicates cases where the model fails to generate a valid cascade at all. The results are for Claude-3.7-Sonnet on PBEBench-Lite, and success denotes Pass@1 = 1 with a sampling budget of 1.
Figure 35: Relation Type Confusion Matrix for Failure (Claude-4 Sonnet (Thinking) PBEBench-Lite): Confusion matrix showing the distribution of relation types in the model prediction vs ground truth. INVALID category indicates cases where the model fails to generate a valid cascade at all. The results are for Claude-4 Sonnet (Thinking) on PBEBench-Lite, and failure denotes Pass@1 = 0 with a sampling budget of 1.
Figure 36: Relation Type Confusion Matrix for Success (Claude-4 Sonnet (Thinking) PBEBench-Lite): Confusion matrix showing the distribution of relation types in the model prediction vs ground truth. INVALID category indicates cases where the model fails to generate a valid cascade at all. The results are for Claude-4 Sonnet (Thinking) on PBEBench-Lite, and success denotes Pass@1 = 1 with a sampling budget of 1.
Figure 37: Relation Type Confusion Matrix for Failure (Claude-4 Sonnet PBEBench-Lite): Confusion matrix showing the distribution of relation types in the model prediction vs ground truth. INVALID category indicates cases where the model fails to generate a valid cascade at all. The results are for Claude-4 Sonnet on PBEBench-Lite, and failure denotes Pass@1 = 0 with a sampling budget of 1.
Figure 38: Relation Type Confusion Matrix for Success (Claude-4 Sonnet PBEBench-Lite): Confusion matrix showing the distribution of relation types in the model prediction vs ground truth. INVALID category indicates cases where the model fails to generate a valid cascade at all. The results are for Claude-4 Sonnet on PBEBench-Lite, and success denotes Pass@1 = 1 with a sampling budget of 1.
Figure 39: Relation Type Confusion Matrix for Failure (Codestral-22B PBEBench-Lite): Confusion matrix showing the distribution of relation types in the model prediction vs ground truth. INVALID category indicates cases where the model fails to generate a valid cascade at all. The results are for Codestral-22B on PBEBench-Lite, and failure denotes Pass@1 = 0 with a sampling budget of 1.
Figure 40: Relation Type Confusion Matrix for Success (Codestral-22B PBEBench-Lite): Confusion matrix showing the distribution of relation types in the model prediction vs ground truth. INVALID category indicates cases where the model fails to generate a valid cascade at all. The results are for Codestral-22B on PBEBench-Lite, and success denotes Pass@1 = 1 with a sampling budget of 1.
Figure 41: Relation Type Confusion Matrix for Failure (DeepSeek-R1-Distill-Qwen-32B PBEBench-Lite): Confusion matrix showing the distribution of relation types in the model prediction vs ground truth. INVALID category indicates cases where the model fails to generate a valid cascade at all. The results are for DeepSeek-R1-Distill-Qwen-32B on PBEBench-Lite, and failure denotes Pass@1 = 0 with a sampling budget of 1.
Figure 42: Relation Type Confusion Matrix for Success (DeepSeek-R1-Distill-Qwen-32B PBEBench-Lite): Confusion matrix showing the distribution of relation types in the model prediction vs ground truth. INVALID category indicates cases where the model fails to generate a valid cascade at all. The results are for DeepSeek-R1-Distill-Qwen-32B on PBEBench-Lite, and success denotes Pass@1 = 1 with a sampling budget of 1.
Figure 43: Relation Type Confusion Matrix for Failure (Gemini 2.5 Flash PBEBench-Lite): Confusion matrix showing the distribution of relation types in the model prediction vs ground truth. INVALID category indicates cases where the model fails to generate a valid cascade at all. The results are for Gemini 2.5 Flash on PBEBench-Lite, and failure denotes Pass@1 = 0 with a sampling budget of 1.
Figure 44: Relation Type Confusion Matrix for Success (Gemini 2.5 Flash PBEBench-Lite): Confusion matrix showing the distribution of relation types in the model prediction vs ground truth. INVALID category indicates cases where the model fails to generate a valid cascade at all. The results are for Gemini 2.5 Flash on PBEBench-Lite, and success denotes Pass@1 = 1 with a sampling budget of 1.
Figure 45: Relation Type Confusion Matrix for Failure (QwQ-32B PBEBench-Lite): Confusion matrix showing the distribution of relation types in the model prediction vs ground truth. INVALID category indicates cases where the model fails to generate a valid cascade at all. The results are for QwQ-32B on PBEBench-Lite, and failure denotes Pass@1 = 0 with a sampling budget of 1.
Figure 46: Relation Type Confusion Matrix for Success (QwQ-32B PBEBench-Lite): Confusion matrix showing the distribution of relation types in the model prediction vs ground truth. INVALID category indicates cases where the model fails to generate a valid cascade at all. The results are for QwQ-32B on PBEBench-Lite, and success denotes Pass@1 = 1 with a sampling budget of 1.
Figure 47: Relation Type Confusion Matrix for Failure (Qwen2.5-32B-Instruct PBEBench-Lite): Confusion matrix showing the distribution of relation types in the model prediction vs ground truth. INVALID category indicates cases where the model fails to generate a valid cascade at all. The results are for Qwen2.5-32B-Instruct on PBEBench-Lite, and failure denotes Pass@1 = 0 with a sampling budget of 1.
Figure 48: Relation Type Confusion Matrix for Success (Qwen2.5-32B-Instruct PBEBench-Lite): Confusion matrix showing the distribution of relation types in the model prediction vs ground truth. INVALID category indicates cases where the model fails to generate a valid cascade at all. The results are for Qwen2.5-32B-Instruct on PBEBench-Lite, and success denotes Pass@1 = 1 with a sampling budget of 1.
Figure 49: Relation Type Confusion Matrix for Failure (Qwen2.5Coder-32B-Instruct PBEBench-Lite): Confusion matrix showing the distribution of relation types in the model prediction vs ground truth. INVALID category indicates cases where the model fails to generate a valid cascade at all. The results are for Qwen2.5Coder-32B-Instruct on PBEBench-Lite, and failure denotes Pass@1 = 0 with a sampling budget of 1.
Figure 50: Relation Type Confusion Matrix for Success (Qwen2.5Coder-32B-Instruct PBEBench-Lite): Confusion matrix showing the distribution of relation types in the model prediction vs ground truth. INVALID category indicates cases where the model fails to generate a valid cascade at all. The results are for Qwen2.5Coder-32B-Instruct on PBEBench-Lite, and success denotes Pass@1 = 1 with a sampling budget of 1.
Figure 51: Relation Type Confusion Matrix for Failure (Qwen3-30B-A3B PBEBench-Lite): Confusion matrix showing the distribution of relation types in the model prediction vs ground truth. INVALID category indicates cases where the model fails to generate a valid cascade at all. The results are for Qwen3-30B-A3B on PBEBench-Lite, and failure denotes Pass@1 = 0 with a sampling budget of 1.
Figure 52: Relation Type Confusion Matrix for Success (Qwen3-30B-A3B PBEBench-Lite): Confusion matrix showing the distribution of relation types in the model prediction vs ground truth. INVALID category indicates cases where the model fails to generate a valid cascade at all. The results are for Qwen3-30B-A3B on PBEBench-Lite, and success denotes Pass@1 = 1 with a sampling budget of 1.
Figure 53: Relation Type Confusion Matrix for Failure (Qwen3-32B (with CoT) PBEBench-Lite): Confusion matrix showing the distribution of relation types in the model prediction vs ground truth. INVALID category indicates cases where the model fails to generate a valid cascade at all. The results are for Qwen3-32B (with CoT) on PBEBench-Lite, and failure denotes Pass@1 = 0 with a sampling budget of 1.
Figure 54: Relation Type Confusion Matrix for Success (Qwen3-32B (with CoT) PBEBench-Lite): Confusion matrix showing the distribution of relation types in the model prediction vs ground truth. INVALID category indicates cases where the model fails to generate a valid cascade at all. The results are for Qwen3-32B (with CoT) on PBEBench-Lite, and success denotes Pass@1 = 1 with a sampling budget of 1.
Figure 55: Relation Type Confusion Matrix for Failure (Qwen3-32B PBEBench-Lite): Confusion matrix showing the distribution of relation types in the model prediction vs ground truth. INVALID category indicates cases where the model fails to generate a valid cascade at all. The results are for Qwen3-32B on PBEBench-Lite, and failure denotes Pass@1 = 0 with a sampling budget of 1.
Figure 56: Relation Type Confusion Matrix for Success (Qwen3-32B PBEBench-Lite): Confusion matrix showing the distribution of relation types in the model prediction vs ground truth. INVALID category indicates cases where the model fails to generate a valid cascade at all. The results are for Qwen3-32B on PBEBench-Lite, and success denotes Pass@1 = 1 with a sampling budget of 1.
Figure 57: Relation Type Confusion Matrix for Failure (Qwen3-Coder-30B-A3B-Instruct PBEBench-Lite): Confusion matrix showing the distribution of relation types in the model prediction vs ground truth. INVALID category indicates cases where the model fails to generate a valid cascade at all. The results are for Qwen3-Coder-30B-A3B-Instruct on PBEBench-Lite, and failure denotes Pass@1 = 0 with a sampling budget of 1.
Figure 58: Relation Type Confusion Matrix for Success (Qwen3-Coder-30B-A3B-Instruct PBEBench-Lite): Confusion matrix showing the distribution of relation types in the model prediction vs ground truth. INVALID category indicates cases where the model fails to generate a valid cascade at all. The results are for Qwen3-Coder-30B-A3B-Instruct on PBEBench-Lite, and success denotes Pass@1 = 1 with a sampling budget of 1.
Figure 59: Relation Type Confusion Matrix for Failure (o3-mini PBEBench-Lite): Confusion matrix showing the distribution of relation types in the model prediction vs ground truth. INVALID category indicates cases where the model fails to generate a valid cascade at all. The results are for o3-mini on PBEBench-Lite, and failure denotes Pass@1 = 0 with a sampling budget of 1.
Figure 60: Relation Type Confusion Matrix for Success (o3-mini PBEBench-Lite): Confusion matrix showing the distribution of relation types in the model prediction vs ground truth. INVALID category indicates cases where the model fails to generate a valid cascade at all. The results are for o3-mini on PBEBench-Lite, and success denotes Pass@1 = 1 with a sampling budget of 1.
Figure 61: Relation Type Confusion Matrix for Failure (o4-mini PBEBench-Lite): Confusion matrix showing the distribution of relation types in the model prediction vs ground truth. INVALID category indicates cases where the model fails to generate a valid cascade at all. The results are for o4-mini on PBEBench-Lite, and failure denotes Pass@1 = 0 with a sampling budget of 1.
Figure 62: Relation Type Confusion Matrix for Success (o4-mini PBEBench-Lite): Confusion matrix showing the distribution of relation types in the model prediction vs ground truth. INVALID category indicates cases where the model fails to generate a valid cascade at all. The results are for o4-mini on PBEBench-Lite, and success denotes Pass@1 = 1 with a sampling budget of 1.

### v4

@@ F.4 Effect of More Examples
F.5 Real SLI Results
We evaluate open-source LLMs on the real SLI dataset using Pass@1, Edit_Sim, and Valid Rate, with results reported in Table 13. Evaluation is performed on 21 PBE instances under substantially relaxed decoding constraints to estimate an upper bound on model performance. Specifically, we use a maximum sequence length four times larger than the corresponding PBEBench-Lite settings and a sampling budget of 32 attempts per instance. Models are additionally allowed to generate cascades of up to length 50, with each rewrite rule permitting $\alpha$ and $\beta$ substrings in the replace($\alpha$, $\beta$) of up to five characters. Despite these concessions, all evaluated models struggle on real SLI. Only the gpt-oss family, and DeepSeek-R1-Distill-Qwen and Qwen3-32B with chain-of-thought prompting achieve Pass@1 of 10% or higher. Among them, gpt-oss-120b performs the best, reaching a Pass@1 of 33%.
Model Max Seq Len First Code Block Last Code Block
Pass@1 Edit Sim Valid Rate Pass@1 Edit Sim Valid Rate
Codestral-22B 8192 0 31 81 0 31 81
Qwen2.5-32B-Instruct 8192 0 32 88 0 32 88
Qwen2.5Coder-32B-Instruct 8192 0 31 82 0 33 80
QwQ-32B 32768 5 63 94 5 63 94
Qwen/Qwen3-32B (with CoT) 32768 10 60 97 10 60 97
Qwen/Qwen3-32B 8192 5 60 79 5 60 79
Qwen3-30B-A3B 32768 5 9 97 5 9 97
DeepSeek-R1-Distill-Qwen-32B 32768 10 51 95 10 51 95
gpt-oss-20B (high) 32768 24 82 88 24 82 88
gpt-oss-120B (high) 32768 33 86 88 33 86 88
Qwen/Qwen3-Coder-30B-A3B-Instruct 8192 5 23 87 5 20 96
Table 13: Real SLI Performance: We evaluate all open source LLMs using the same metrics and prompts as PBEBench-Lite. Each LLM is given 32 attempts to solve the 21 PBE instances and each instance contains 50 input-output examples.
Dataset r p
CLUTRR 0.6364 0.035287
SLR-Bench 0.8091 0.002559
PBEBench-Lite 0.8273 0.001677
Table 14: Benchmark Ranking Correlations: Spearman rank correlations between open-source model rankings induced by inductive reasoning benchmarks and real SLI performance. For PBEBench-Lite and real SLI, rankings we use Pass@1 as the primary metric with Edit_Sim as a tie breaker. For CLUTRR, Acc and NNAcc are used, while for SLR-Bench, Acc and Partial Score serve as the primary and tie-breaking metrics, respectively.
F.6 Efficiency of Problem Proposer
@@
Program Ordering Prompt You are solving a program ordering puzzle. Given input-output string pairs and a scrambled list of string replacement programs, your goal is to determine the correct execution order.
## Background
Each program performs a Python string replacement: replace("A", "B") replaces all occurrences of "A" with "B". Why order matters: Feeding: One program creates substrings that another program can match.
Example: replace("a","bc") followed by replace("bc","x").
Bleeding: One program removes substrings that another program would have matched.
Example: replace("ab","x") followed by replace("a","y").
## Your Task
**Inputs:** {inputs}
**Outputs:** {outputs}
**Scrambled Programs** (indices 0 to {n_minus_1}):
{programs_formatted}
[…]
## Approach
1. Trace through what each program does 2. Identify potential feeding/bleeding interactions 3. Reason about which programs must come before others 4. Verify your ordering produces the expected outputs
## Output Format
Provide your final answer as a JSON array of indices:
‘‘‘json
[i0, i1, i2, ...]
Your ordering must be a permutation of [0, 1, ..., {n_minus_1}].
F.7 Factorial Analysis
@@
F.7.1 Effect of Long Chain of Thought Reasoning
@@
F.7.2 Analysis of Models of Varying Capabilities
@@
Term Coefficient Interpretation
cascade_length -0.521
Strongest predictor.
Longer = harder.
counterfeeding -0.291 Hardest relation.
counterbleeding -0.068 Moderate difficulty.
bleeding -0.034
Small negative
effect.
feeding -0.004 Negligible effect.
Table 17: Difficulty Analysis of PBEBench-Lite-Perm: Logistic regression coefficients for factors affecting the difficulty of program reordering and interpretation of the values.
@@
F.8 Logistic Regression Analysis
@@
Cascade Length Pass@1
@@
Sampling Budget First Code Block Last Code Block Nulls
@@
Sampling Budget First Code Block Last Code Block Nulls
@@
Max Seq Length First Code Block Last Code Block Nulls
@@
F.9 Qualitative Analysis of Program Reordering Performance
We present a qualitative analysis of reasoning strategies employed by large language models on a program reordering task grounded in phonological rule ordering. Through examination of model responses across systems with visible chain-of-thought, we identify two distinct reasoning paradigms: (1) template-based constraint reasoning that explicitly models feeding/bleeding interactions between rules, and (2) exhaustive enumeration via brute-force permutation testing with forward simulation. Models employing constraint-based reasoning (e.g., Codestral-22B) correctly identify feeding relationships but often derive incorrect ordering constraints—stating “Program A feeds Program B” but concluding A must precede B, the opposite of correct inference. Meanwhile, enumeration-based approaches (e.g., QwQ-32B, DeepSeek-R1) exhibit combinatorial collapse as problem size increases: response length grows from roughly 7,000 to about 28,000 characters while accuracy drops from nearly 90% to around 40% as the search space expands from 2 to 120 permutations. Notably, models rarely reference counterfeeding or counterbleeding despite these interactions being critical to the hardest instances. These findings reveal systematic gaps in how LLMs translate domain knowledge into valid constraint satisfaction and highlight the limitations of brute-force search for combinatorial reasoning tasks.
F.10 Confusion Matrices for Cascade Lengths
@@
F.11 Confusion Matrices for BFCC Relations
@@
Appendix G Use of AI Assistants
We made limited use of AI assistants during the preparation of this manuscript, primarily for rephrasing, paraphrasing, and performing grammatical and stylistic checks. All technical content, analysis, and conclusions are solely those of the authors.